# Closing the Gap: Achieving Better Accuracy-Robustness Tradeoffs Against Query-Based Attacks

### 15 Dec 2023 v1

@@ Conclusion
In this paper, we showed that limiting the invocation of a test-time defense to low-confidence inputs might be sufficient to obstruct the search for adversarial directions in query-based attacks. Our approach can be applied generically to existing test-time defenses and is training-free. We therefore hope that our findings motivate further research in this area.
@@ Acknowledgments
We provide the Pareto plot in Figure 9 for RND and RCR, where we see an improvement in $\mathsf{CA}/\mathsf{RA}$ of up to $1\%$ over the baseline. We also notice that the $\mathsf{RA}$ decreases with increasing the parameters $\nu$ and $\tau$. This is related to the vastly larger input space and the higher number of classes, for which introduced noise can result in easier label flips, which actually boosts the success of an adversary in the untargeted attack setting.
(a) ImageNet, $\mathsf{D}$ = RND
(b) ImageNet, $\mathsf{D}$ = RCR
Figure 9: Pareto frontier when applying our approach in conjunction with RND and RCR against PSJA on ImageNet.
@@
dheu6P33fBwB4ngcAcByHJpPJl+fn54mD3Gg0NrquXxeLRQAAwzAYj8cwTZPwPH9/sVg8PXweDAauqqr2cDjEer1GJBLBZDJBs9mE4zjwfZ85lAGg2+06hmGgXq+j3+/DsixYlgVN03a9Xu8jgCNCyIegIAgx13Vfd7vdu+FweG8YRkjXdWy329+dTgeSJD3ieZ7RNO0VAXAPwDEAO5VKndi2fWrb9jWl9Esul6PZbDY9Go1OZ7PZ9z/lyuD3OozU2wAAAABJRU5ErkJggg==" alt="[LOGO]">

### 21 Mar 2024 v2

@@ Conclusion
In this paper, we showed that limiting the invocation of an inference-time defense to low-confidence inputs might be sufficient to obstruct the search for adversarial samples in query-based attacks. Our approach can be applied generically to existing inference-time defenses and is training-free. We therefore hope to motivate further research in this area.
@@ Acknowledgments
We notice that the $\mathsf{RA}$ decreases with increasing the parameters $\nu$ and $\tau$. This is related to the vastly larger input space and the higher number of classes, for which introduced noise can result in easier label flips, which actually boosts the success of an adversary in the untargeted attack setting.